Ions B and C react to formthe complex BC. If 35.0 mL of 1.00MB is combined with 35.0mL of 1.00M
C, 0.00500 mol of BC is formed. Determine the equilibrium constant for this reaction.
1
Expert's answer
2020-09-17T07:15:00-0400

After mixing together: B + C = BC

volume is 70 ml

mole of BC is 0.00500

molea of B and C are equal 0.035*1 - 0.005 = 0.030

concentrations are [BC] = 0.005/0.070 = 0.0714 M,

[B] = [C] = 0.030/0.070 = 0.4286 M

Equilibrium constant is

K = [BC]/([B]*[C]) = 0.3887